If current can flow through a capacitor supplied with dc to complete the circuit then how can we say that a capacitor can block dc??


 

Grade:11

3 Answers

bkarthik iyer
12 years ago

2011-09-25

 

b'coz xc=1/2∏fC

 

but for dc f=0

thus xc=∞

thus I=0

that is why it blocks current.

Ashwin Muralidharan IIT Madras
12 years ago

Hi Deeksha,

 

Please note that a capacitor blocks DC only when it is fully charged.

 

So you apply a potential of V across a capaitor of cpacitance C. Then if it is charged to a value Q = CV, in that case it would block a direct current.

Even when it blocks a direct current, please note that it has a potential difference maintained across its ends, due to its charge, given by V = Q/C.

 

As long as this charge is there on the capacitor, it would block DC.

Note that current I = dq/dt = d/dt(CV). As long as a constant voltage is applied, d/dt(CV) = 0.

Hence we say that, it blocks the DC.

 

Hope that was helpful.

 

All the best,

Regards,

Ashwin (IIT Madras).

deeksha sharma
12 years ago

But why is it that while finding out the expression for displacement current we assume that the circuit is completed when a capacitor is connected in series with a bulb